OpenStudy (anonymous):

A right triangle has base (x - 7) units and height (2x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work.

OpenStudy (anonymous):

@Hero

hero (hero):

 In general when you have a right triangle, a and b are considered the side lengths and c is considered the hypotenuse. The pythagorean theorem is an equation that relates all three sides together: a^2 + b^2 = c^2 Where c^2 is the square length of the hypotenuse. In this case, we can say that a = the height of the right triangle = 2x - 10 b = the base of the right triangle = x - 7 Therefore the square length of the hypotenuse is represented by: (2x - 10)^2 + (x - 7)^2 = c^2 Basically what you want to do is expand the left side completely to find the appropriate expression for c^2